On the paper there is a right triangle it has the two lines on two of the sides to mean they are the same. It doesn't tell any of the lengths but if is on a grid. It has 4 squares on both of the sides that are the same, however I'm not sure if it's drawn to scale. It wants me to find the slope and use the Pythagorean theorem if necessary.
Answer by Theo(13342)

if it's on a grid and each of the equal legs have 4 squares each, then you can assume that each is 4 units in length, whatever the unit is that they are measuring with.
if the triangle is slanting up toward the right, then the slope will be positive.

if the triangle is slanting down toward the right, then the slope will be negative.

the slope is the opposite side divided by the adjacent side which would be 4/4 = 1 if slanting up and equal -1 if slanting down.

if you want to find the length of the hypotenuse then you need to use the pythagorean formula of a^2 + b^2 = c^2.

you will get 4^2 + 4^2 = 16 + 16 = 32 which is the square of the hypotenuse.

the hypotenuse length becomes the square root of 32 which can be simplified to:

4 times (the square root of 2) units.